For the given functions f and g, find the indicated composition.

f(x) = 4x2 + 2x + 7, g(x) = 2x - 3
(g∘f)(x)

Answer by ewatrrr(24785)


Hi

f(x) = 4x^2 + 2x + 7,	g(x) = 2x - 3
(g∘f)(x) = 2(4x^2 + 2x + 7) - 3 = 8x^2 + 4x + 11
Wish You the Best in your Studies.